2. For the following reactions, write the balanced molecular equation, the complete ionic equation and the net ionic equation. Make sure that the equations are balanced. You should write three equations. (NH4)3P04(aq) + Zn(NO3)2 (aq) → NH4NO3(aq) + Zn3(POJ2(s) OH)2
media%2F329%2F32935a7c-4c44-4b70-88d9-e5
Answer #1

Answer:

the balanced equations are

2 (NH4)3PO4(aq) + 3 Zn(NO3)2(aq) = 6 NH4NO3(aq) + Zn3(PO4)2(s)

3 Zn(NO3)2 + 2 (NH4)3PO4 = Zn3(PO4)2 (s) + 6 NH4NO3 (aq)

3 Zn2+ + 6 NO3- + 6 NH4+ + 2 PO43- = Zn3(PO4)2 + 6 NH4+ + 6 NO3-

3 Zn2+ + 2 PO43- = Zn3(PO4)2

Na2CO3 + 2 HCl = 2 NaCl + CO2 + H2O

2Na+ (aq) + CO3 2- (aq) + 2H+ (aq) + 2Cl-(aq) --------> CO2 (g) + H2O (l) + 2Na+ (aq) + 2Cl- (aq)

CO3 2- (aq) + 2H+ (aq) --------> CO2 (g) + H2O (l)

Thanks :)